The area of a triangle, in cm2, whose sides have length 6 cm, 7 cm, and 9 cm is:
A 2√110 b 12√3 c 9√6 d 22 e none of these

Answer by Greenfinch(383)

Using Area = sqrt (s)(s-a)(s-b)(s-c)
s is 1/2 (6+7+9)= 11
Gives sqrt(11 x 5 x 4 x 2) = sqrt 440 = 2sqrt 110
So answer is A
Alternatively use cosine rule
cos A= (6^2 + 7^2 - 9^2)/2 x 6 x 7 = 4/84 = 1/21
A = 87.27 degrees
so sin A = 1 - (1/21)^2 = 440/441
Area = 1/2 (6 x 7) sin A = 21 x 440/441 = 20.95
(2 sqrt 110 = 20.98)